OpenStudy (anonymous):

Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and 3 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Keven needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? 3 feet 6 feet 8 feet 10 feet I got 6 feet for my answer, but I'm not sure if I did this right...Could someone check this for me?

OpenStudy (campbell_st):

start by finding the volume of concrete he bought. 72.59/98 = then convert the known measurements to yards 10ft = 10/3 ft 3 inches = 0.25/3 and I don't think its 6 ft

OpenStudy (anonymous):

Okay, so I did: 0.74 = 3.3 * 0.083 * w 0.74 = 0.27 * w 0.74/0.27 = 2.7 w = 2.7 So the answer would be 3 feet?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Thank you for your help (: